OpenStudy (anonymous):

solve for y: x=(3y-2)/(2y+3)

OpenStudy (anonymous):

2xy + 3x = 3y - 2 3x + 2 = 3y - 2xy 3x + 2 = y (3 - 2x) hence y = (3x + 2)/ (3 - 2x)